PART A Hardcopy due 11 am Moody December 12 A m, 5


how do I factor in collision and friction
Answer #1

Consider the motion on the rough track

Wf = work done by friction = 5 J

Vi = initial speed at the top = 20 cm /s = 0.2 m/s
Vf = final speed at the bottom = ?

h = height of track = 3 m

using conservation of energy between Top and bottom track

(0.5) m Vi2 + mg h = (0.5) m Vf2 + Wf

(0.5) (0.5) (0.2)2 + (0.5) (9.8) (3) = (0.5) (0.5) Vf2 + 5

Vf = 6.23 m/s

V = velocity of combination after collision

M = mass of block = 1 kg

using conservation of momentum

m Vf = (m + M) V

(0.5) (6.23) = (0.5 + 1) V

V = 2.08 m/s

b)

KE lost = (0.5) m Vf2 - (0.5) (m + M) V2 = (0.5) (0.5) (6.23)2 - (0.5) (0.5 + 1) (2.08)2 = 6.5 J

c)

consider the vertical motion of the coupled mass

Voy = initial velocity = 0 m/s

a = acceleration = - 9.8

t = time

Y = vertical displacement = - 2m

using the equation

Y = Voy t + (0.5) a t2

- 2 = 0 t + (0.5) (-9.8) t2

t = 0.64 sec

along the horizontal direction

Vox = initial velocity along horizontal direction =

X = Vox t = 2.08 x 0.64 = 1.33 m